OpenStudy (anonymous):

three identical dice are rolled the probability that same number will appear on each of them is?

OpenStudy (sidsiddhartha):

if three dices are rolled then total number of cases=6*6*6 idendical numbers can be =(1,1,1) (2,2,2) (3,3,3) (4,4,4) (5,5,5) (6,6,6) like this so here 6 cases are available so ur required probability=\[\frac{ 6 }{ 6*6*6 }=1/36\] ok!!!